David has 40 coins in his bank. If the coins consists of quarters and dimes and value $7.60, how many of each coin does he have?
Answer by kiru_khandelwal(79)

Let the
number of quarters = x
number of dimes = y
Since david has 40 coins, so
x+y = 40 ...........Equation (1)
1 quarter = 25 cents...so x quarters = 25*x cents
1 dime = 10 cents.....so y dimes = 10*y cents
$7.60 = 7.60 * 100 = 760 cents
Since the total value is $7.60 or 760 cents
Therefore,
25x+10y = 760.........Equation (2)
Solving equation (2)
5*5x + 5*2y = 5*152
=> 5(5x+2y) = 5(152)
=> 5x+2y=152.........Equation (3)
From Equation(1)
y = 40-x
Substituting the value of y in Equation(3)
5x + 2y = 152
=> 5x + 2(40-x) = 152
=> 5x + 80 - 2x = 152
=> 3x = 152-80
=> 3x = 72
=> x = 24
Therefore, No. of Quarters = 24
As, y = 40 -x
=> y = 40-24 = 16
Number of dimes = 16
Hence solved